# generalised polynomial chaos

### v1

@@ I Introduction
distributed energy resources are increasingly participating in ancillary service (AS) markets. Many of these resources. such as heat pumps, electric vehicles and photovoltaic (PV) installations have uncertain capabilities. To facilitate the participation of these DERs, some transmission system operators already allow bids for which the full power availability is only guaranteed with a certain probability [1]. This has important consequences for the reliable procurement of reserve capacity. Add short review on approaches for reserve procurement. Look for probabilistic approaches + especially considering dependencies of multiple DER types.
@@
Recently, generalised polynomial chaos (gPC) has been used in power systems applications to model propagation of uncertain input variables. Advantages include the ability to model general probability distributions and represent stochastic variables through orthogonal polynomial expansions. This approach was first used for stochastic power flow computations [4], with inequality constraints formulated only with respect to the expected value. Later work, such as [5], introduces the use of gPC for the solution of optimal power flow problems. Typically, approaches modelling uncertainty with gPC formulate chance constraints through a moment-based reformulation. For Gaussian random variables or distributionally robust (DR) uncertainty, exact reformulations are found [6]. For other distributions, [3] proposes an iterative tuning approach for the bounds in the moment-based reformulation in an attempt to recover the desired constraint satisfaction rate. A major limitation of these approaches, is that they consider the different random input variables to be independent. To allow modelling dependencies between random variables, [7] proposes to construct an orthonormal polynomial basis using a Gram-Schmidt (GS) procedure. A challenge with this approach is that it leads to numerical issues due to the poor conditioning of the GS procedure[2]. Other methods to model dependent random variables with gPC exist, as reported for example in [2]. The Rosenblatt transformation is a mapping that transforms dependent random variables to a space of independent random variables, in which an orthonormal polynomial basis can be constructed using standard approaches. This has for example been applied to solve stochastic power flows [9]. However, this approach requires to determine the inverse Rosenblatt transform for each sample used to obtain polynomial chaos expansion (PCE) coefficients. here further works on dependent gPC?
@@
None of the works above allow to assess the probability of violating the power reserve requirements when considering uncertain bids exhibiting dependencies. Additionally, the proposed approaches to extend gPC to dependent variables, lead to computational issues due to error propagation and computation time. Therefore, this paper makes the following contributions:
@@ II Methodology
Consider the random vector $\bm{\xi}=[\xi_{1},\ldots,\xi_{d}]^{T}$ with joint PDF $p(\bm{\xi})$ and support $\Xi$.
A known method to deal with uncertainty (is the use of a Polynomial Chaos Expansion (PCE), casting a function $f(\bm{\xi})$ of random variables $\bm{\xi}$ into a sum of orthonormal polynomials $\psi(\bm{\xi})$:
$f(\bm{\xi})=\sum_{l=0}^{\infty}a_{l}\psi_{l}(\bm{\xi}),$ (1)
where orthonormality is required with respect to the expectation value
$\mathbb{E}[\psi_{m}(\bm{\xi})\psi_{n}(\bm{\xi})]=\int_{\Xi}\psi_{m}(\bm{\xi})\psi_{n}(\bm{\xi})p(\bm{\xi})d\bm{\xi}=\delta_{mn}.$ (2)
The coefficients can be computed by projection onto the polynomial basis, where the scalar product is identical to the expectation value of the product:
$a_{l}=\left\langle f(\bm{\xi}),\psi_{l}(\bm{\xi})\right\rangle=\mathbb{E}[f(\bm{\xi})\psi_{l}(\bm{\xi})]$ (3)
The main difficulty is the search for orthonormal polynomials with respect to the arbitrary PDF $p(\xi)$. As proven in (XXXSomething), polynomials arising from Askey-scheme are orthogonal for their respective probability distribution (e.g. Hermite polynomials for Gaussian PDFs), however, this assumes independence between the elements of the random vector. In turn, Navarro et al. suggested in [7] that for arbitrarily correlated random variables, orthonormal polynomials can be constructed from a set of linearly independent polynomials using the Gram-Schmidt orthonormalization scheme.
@@ II-B Joint Probability using Copulas
For practical applications, the joint PDF $p(\bm{\xi})$ needs to be approximated in a tractable manner. This requires both a model of the marginal PDFs of each random variable $\xi_{i}$ as well as their dependence. Dependence modelling can be achieved by using a copula $C$, a function mapping the cumulative distribution functions (CDF) $F_{i}(\xi_{i})$ of each marginal to a joint CDF $F(\bm{\xi})$:
$F(\bm{\xi})=C(F_{1}(\xi_{1}),\ldots,F_{d}(\xi_{d}))$ (4)
As the copula takes CDFs as inputs, it effectively decouples the marginal distributions from the dependence structure. According to Sklar’s theorem (XXX), there exists a unique copula for every possible joint CDF. In practice, one can use an empirical copula, as for example in (Pål paper?), which makes no assumptions on the correlation structure or fit a parametric copula. Archimedean copulas such as Gumbel or Clayton copulas, have a single parameter to capture upper and lower tail-dependence respectively. Gaussian, or more generally Student-t copulas, are defined by a correlation matrix $\Sigma$ and degrees of freedom $\nu$, where in the limit $\nu\rightarrow\infty$, the Student-t becomes a Gaussian copula.
For the application reported here, featuring a high dimensionality a Gaussian copula was chosen. This allows for scalability and mathematically convenient expressions (XXX and what?). Setting $F_{i}(\xi_{i})=u_{i}$, it is defined as:
$C_{\bm{\Sigma}}(u_{1},...,u_{d})=\Phi_{\bm{\Sigma}}(\Phi^{-1}(u_{1}),...,\Phi^{-1}(u_{d}))$ (5)
where $\Phi^{-1}(u_{i})$ is the inverse of the normal CDF and $\Phi_{\bm{\Sigma}}(\bm{z})$ the CDF of the multivariate normal with correlation matrix $\bm{\Sigma}$. The joint PDF of the random vector $\bm{\xi}$ is then expressed as
@@
where $c(\bm{u})$ is the copula density
$c(\bm{u})=\frac{\partial^{d}C(u_{1},\dots,u_{d})}{\partial u_{1}\,\partial u_{2}\cdots\partial u_{d}}.$ (7)
In the case of a Gaussian copula, this results in
$c_{\Sigma}(\bm{u})=\frac{\phi_{\bm{\Sigma}}\big(\Phi^{-1}(u_{1}),\ldots,\Phi^{-1}(u_{d})\big)}{\prod_{i=1}^{d}\phi\big(\Phi^{-1}(u_{i})\big)},$ (8)
where $\phi_{\bm{\Sigma}}(\bm{x})$ is the PDF of the multivariate normal with correlation matrix $\bm{\Sigma}$} and $\phi(x_{i})$ the normal PDF.
@@ II-C Quadrature Integration
Both for the computation of the orthonormal polynomials using GS and for the projection of expansion coefficients in Eq. (3), integrals of the form
@@
have to be evaluated numerically, where $g(\bm{\xi})$ is an arbitrary function. Given the high dimensionality of the integral, evaluation with Monte Carlo sampling or importance sampling schemes would be standard, this however leads to relatively high errors. Further, importance sampling schemes are difficult to implement, as the integrands do not follow a constant distribution
To allow accurate evaluation, quadrature integration is used, this however suffers from computational tractability problems if applied to high-dimensional problems. For this reason a scheme reducing the integration dimension is introduced. First, the integration domain needs to be transformed to a space with known quadrature rules. One such space is the unit cube $0,1]^{d}$ over which the Gauss-Legendre rules can be applied. thus let
@@
The Jacobian matrix of this transformation is diagonal with entries $\partial u_{i}/\partial\xi_{i}=f_{i}(\xi_{i})$, hence
@@
Applying the change of variables and substituting in the definition of $p(\bm{\xi})$ from Eq. (6) yields
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{[0,1]^{d}}f\big(\bm{F}^{-1}(\bm{u})\big)\,c(\bm{u})\,d\bm{u}$ (12)
where $\bm{F}^{-1}(\bm{u})$ is the componentwise inverse CDF. This expression makes no assumptions on the copula used and hence… However, as the copula density $c(\bm{u})$ is in general neither polynomial nor constant, the quadrature will be inexact. In the specific case of Gaussian copulas, a transformation to latent Gaussian space
$z_{i}=\Phi^{-1}(u_{i})\in(-\infty,\infty),\qquad i=1,\dots,d.$ (13)
is helpful. The Jacobian matrix is again diagonal with entries $\partial u_{i}/\partial z_{i}=\phi(z_{i})$, hence
$d\bm{u}=\Big(\prod_{i=1}^{d}\phi(z_{i})\Big)\,d\bm{z}.$ (14)
Applying the change of variables to the copula density and using the definition of the Gaussian copula density from Eq. (8) yields
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{\mathbb{R}^{d}}g\big(\bm{F}^{-1}(\Phi(\bm{z}))\big)\;\phi_{\Sigma}(\bm{z})\,d\bm{z}.$ (15)
This form allows for Gauss-Hermite quadrature, which expects a polynomial weighted by an uncorrelated normal distribution. Computing the Cholesky factorisation of the correlation matrix $\bm{\Sigma}=\bm{L}\bm{L^{T}}$ allows to “colour“ the correlation onto the quadrature. Letting $\bm{y}_{n}$ and $w_{n}$ be the quadrature nodes and weights. The linear transformation
$\bm{z_{n}}=\bm{L}\bm{y_{n}}$ (16)
induces the correlation. Overall,
$\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})\,d\bm{x}\;\approx\;\sum_{n=1}^{N}w_{n}\cdot g(\bm{F}^{-1}(\Phi(\bm{Ly_{n}})))$ (17)
Standard Gauss-Hermite quadrature uses a tensor-product grid, meaning evaluating $k$ points per dimension for a $d_{I}$ dimensional integral leads to $N=k^{d_{I}}$ nodes. The exponential scaling with the integral dimension requires a strict limitation.
@@ II-C1 Numerical Improvements
Two ideas allow for decent computational performance when computing the gPC basis. First, GS sequentially evaluates inner products of polynomials to construct the basis polynomials. As noted in [8], the inner product of two arbitrary polynomials $P_{m}(\bm{\xi}),P_{n}(\bm{\xi})$ constructed from the multivariate monomial set $\{1,\xi_{1},\ldots,\xi_{d}^{\nu}\}$, can be expanded to a sum of monomial expectation values
$\left\langle P_{m}(\bm{\xi}),P_{n}(\bm{\xi})\right\rangle$ $=p_{0}\left\langle 1,1\right\rangle+p_{1}\left\langle 1,\xi_{1}\right\rangle+\ldots+p_{P}\left\langle\xi_{d}^{\nu},\xi_{d}^{\nu}\right\rangle$
@@
where $p_{i}$ and $e_{i}$ are the appropriate coefficients. Evaluating the monomial expectation values before applying GS can be done in a parallel fashion, allowing for increased speed.
Second, the expectation values of monomials only depends on dimensions appearing in the monomial, for instance
@@
For elliptic copulas the reduced PDFs have identical structure to the full joint PDF, so Eq. (6) still holds, but with less input variables.
These two observations limit the maximum integral dimension to be evaluated to $2\nu$ with expectations of the form $\mathbb{E}[\xi_{1}\xi_{2}\ldots\xi_{2\nu}]$. Crucially, this decouples integral dimension $d_{I}$ from random variable dimension $d$, avoiding the curse of dimensionality.

### v2

@@ I Introduction
distributed energy resources are increasingly participating in ancillary service (AS) markets. Many of these resources. such as heat pumps, electric vehicles and photovoltaic (PV) installations have uncertain capabilities. To facilitate the participation of these DERs, some transmission system operators already allow bids for which the full power availability is only guaranteed with a certain probability [1]. This has important consequences for the reliable procurement of reserve capacity. This raises new questions with respect to the amount of reserve that should be procured, as the availability of the procured reserve becomes a stochastic quantity.
@@
Recently, generalised polynomial chaos (gPC) has been used in power systems applications to model propagation of uncertain input variables. Advantages include the ability to model general probability distributions and represent stochastic variables through orthogonal polynomial expansions. This approach was first used for stochastic power flow computations [5], with inequality constraints formulated only with respect to the expected value. Later work, such as [6], introduces the use of gPC for the solution of optimal power flow problems. Typically, approaches modelling uncertainty with gPC formulate chance constraints through a moment-based reformulation. For Gaussian random variables or distributionally robust (DR) uncertainty, exact reformulations are found [7]. For other distributions, [3] proposes an iterative tuning approach for the bounds in the moment-based reformulation in an attempt to recover the desired constraint satisfaction rate. A major limitation of these approaches, is that they consider the different random input variables to be independent. To allow modelling dependencies between random variables, [8] proposes to construct an orthonormal polynomial basis using a Gram-Schmidt (GS) procedure. A challenge with this approach is that it leads to numerical issues due to the poor conditioning of the GS procedure[2]. Other methods to model dependent random variables with gPC exist, as reported for example in [2]. The Rosenblatt transformation is a mapping that transforms dependent random variables to a space of independent random variables, in which an orthonormal polynomial basis can be constructed using standard approaches. This has for example been applied to solve stochastic power flows [11]. However, this approach requires to determine the inverse Rosenblatt transform for each sample used to obtain polynomial chaos expansion (PCE) coefficients. For linear dependencies, this can be simplified to the Nataf transform [4], but the authors of [2] show that it performs worse than their proposed GS procedure. Other approaches, so-called domination methods, approximate functions of dependent random variables with functions of independent random variables, but only show good performance when an accurate approximation can be obtained [2].
@@
None of the works above handling reserve procurement allow to assess the probability of violating the power reserve requirements when considering uncertain bids exhibiting dependencies. Additionally, the proposed approaches to extend gPC to dependent variables, lead to computational issues due to error propagation and computation time. Therefore, this paper makes the following contributions:
@@ II Methodology
Consider the random vector $\bm{\xi}=[\xi_{1},\ldots,\xi_{d}]^{\top}$ of dimension $d$ with joint probability density function (PDF) $p(\bm{\xi})$ and support $\Xi$. As shown in previous works (e.g. [8]), a square-integrable random function $q(\bm{\xi})$ can be represented by a PCE
$q(\bm{\xi})\cong\sum_{l=0}^{L}a_{l}\psi_{l}(\bm{\xi}).$ (1)
where the equality becomes exact as the arbitrary number of terms $L$ goes to infinity. The basis polynomials $\{\psi_{l}\}$ are orthonormal with respect to the inner product
$\left\langle\psi_{m},\psi_{n}\right\rangle=\int_{\Xi}\psi_{m}(\bm{\xi})\,\psi_{n}(\bm{\xi})\,p(\bm{\xi})\,\mathrm{d}\bm{\xi}.$ (2)
The expansion coefficients $\{a_{l}\}$ are obtained by projection onto the polynomial basis. Since this inner product corresponds to an expectation, the coefficients read
$a_{l}=\left\langle q,\psi_{l}\right\rangle=\mathbb{E}\!\left[q(\bm{\xi})\,\psi_{l}(\bm{\xi})\right].$ (3)
The main challenge lies in the construction of orthonormal polynomial bases with respect to an arbitrary joint PDF $p(\bm{\xi})$. For independent random variables, orthogonal polynomial families associated with the Askey scheme can be employed, since they are orthogonal with respect to their corresponding marginal distributions (e.g. Hermite polynomials for Gaussian random variables).
For general dependent random variables, the basis can be constructed from a suitably chosen set of multivariate monomials of the random variables (e.g. all monomials with degree $\leq\nu$). As noted, a GS procedure can be used for the construction, but is numerically ill-conditioned [2]. To address this, [10] provides an alternative algorithm based on a Whitening transformation, relying only in part on GS procedures.
Finally, the moments of the output distribution of $q(\bm{\xi})$ can be recovered from the expansion coefficients. For instance the mean and variance are given by
$\mu$ $=a_{0}$ (4)
$\sigma$ $\cong\left\|[a_{1},\ldots,a_{L}]^{\top}\right\|_{2}$
@@ II-B Joint Probability using Copulas
For practical modelling of statistically dependent random variables, the joint PDF $p(\bm{\xi})$ must be approximated in a tractable manner. This requires specifying both the marginal distributions of the individual components $f(\xi_{i})$ as well as their dependence structure.
Copulas provide a general and flexible framework for separating the modelling of marginal distributions from the modelling of statistical dependence. Specifically, a Copula $C$ links the marginal cumulative distribution functions (CDFs) $F_{i}(\xi_{i})$ to the joint CDF $F(\bm{\xi})$. This link is implicitly expressed as
$F(\bm{\xi})=C\!\left(F_{1}(\xi_{1}),\ldots,F_{d}(\xi_{d})\right).$ (5)
Sklar’s Theorem states, that for every multivariate CDF there exists a unique Copula capturing the general dependence structure given by (5). In practice, one of several parametric Copula families is used.
For the application considered in this paper no specific tail-dependence is expected. As a result, a Gaussian Copula is selected to maintain computational tractability and leverage its closed-form mathematical expressions. Identifying $u_{i}=F_{i}(\xi_{i})$, it is defined as:
$C_{\bm{\Sigma}}(\bm{u})=\Phi_{\bm{\Sigma}}(\Phi^{-1}(u_{1}),...,\Phi^{-1}(u_{d}))$ (6)
where $\Phi_{\bm{\Sigma}}(\bm{z})$ is the multivariate normal CDF with correlation matrix $\bm{\Sigma}$ and $\Phi^{-1}(u_{i})$ the inverse normal CDF. The joint PDF is then expressed as
@@
where $c(\bm{u})$ is the Copula density. For a Gaussian Copula, the density is given by
$c_{\bm{\Sigma}}(\bm{u})=\frac{\phi_{\bm{\Sigma}}\big(\Phi^{-1}(u_{1}),\ldots,\Phi^{-1}(u_{d})\big)}{\prod_{i=1}^{d}\phi\big(\Phi^{-1}(u_{i})\big)},$ (8)
where $\phi_{\bm{\Sigma}}(\bm{x})$ is the multivariate normal PDF with correlation matrix $\bm{\Sigma}$} and $\phi(x_{i})$ the normal PDF.
@@ II-C Quadrature Integration
Both the construction of the orthonormal polynomial basis and the computation of the expansion coefficients require the evaluation of a large number of integrals of the form
@@
with $g(\bm{\xi})$ an arbitrary function. In general, these integrals cannot be evaluated analytically and must be approximated numerically. Given the potentially high dimensionality of the integrals, Monte Carlo sampling is a natural choice. This however requires a large amount of samples to achieve sufficient accuracy and sampling errors may accumulate throughout the PCE process. Moreover, the use of importance sampling is not straightforward in this context, since the integrands encountered vary across polynomial orders and random functions.
Therefore, to enable fast and accurate numerical integration, quadrature integration schemes in combination with dimensionality-reduction strategies are used. First, the integration domain is mapped to a space with known quadrature rules. One such space is the unit cube $0,1]^{d}$ on which tensor-grid Gauss-Legendre quadrature can be applied. Let
@@
The Jacobian matrix of this transformation is diagonal with entries $\partial u_{i}/\partial\xi_{i}=f_{i}(\xi_{i})$ the marginal PDFs. Hence
@@
Applying the change of variables and substituting $p(\bm{\xi})$ from (7) yields
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{[0,1]^{d}}g\big(\bm{F}^{-1}(\bm{u})\big)\,c(\bm{u})\,d\bm{u}$ (12)
where $\bm{F}^{-1}(\bm{u})$ is short-hand notation for the componentwise inverse marginal CDF. Evidently, this transformation is valid no matter the Copula used, in principle allowing any Copula family to be used for this quadrature integration scheme. However, the Copula density $c(\bm{u})$ is in general neither polynomial nor constant, meaning Gauss-Legendre quadrature will be inexact.
For the specific case of Gaussian Copulas, consider a second transformation to a Gaussian latent space
$z_{i}=\Phi^{-1}(u_{i})\in(-\infty,\infty),\qquad i=1,\dots,d$ (13)
on which tensor-grid Gauss-Hermite quadrature is applicable. The Jacobian matrix is again diagonal with entries $\partial u_{i}/\partial z_{i}=\phi(z_{i})$, hence
$d\bm{u}=\Big(\prod_{i=1}^{d}\phi(z_{i})\Big)\,d\bm{z}$ (14)
Applying the change of variables and substituting the Gaussian Copula density from (8) yields
$\int_{\Xi}g(\bm{\xi})\cdot p(\bm{\xi})d\bm{\xi}=\int_{\mathbb{R}^{d}}g\big(\bm{F}^{-1}(\bm{\Phi}(\bm{z}))\big)\;\phi_{\bm{\Sigma}}(\bm{z})\,d\bm{z}$ (15)
with $\bm{\Phi}(\bm{z})$ the componentwise normal CDF. The result in (15) features a correlated normal PDF. However, Gauss-Hermite quadrature expects to weighted by an uncorrelated normal PDF. To address this, let $\bm{\Sigma}=\bm{L}\bm{L^{T}}$ be the Cholesky factorisation of the correlation matrix. For a given set of quadrature nodes $\{\bm{z}_{n}\}$ and weights $\{w_{n}\}$ the linear transformation
$\bm{y}_{n}=\bm{L}\bm{z}_{n}$ (16)
allows to “colour“ the correlation onto the quadrature nodes resulting in
$\int_{\Xi}g(\bm{\xi})\,p(\bm{\xi})\,d\bm{x}\;\approx\;\sum_{n=1}^{N}w_{n}\cdot g(\bm{F}^{-1}(\Phi(\bm{Lz}_{n})))$ (17)
The number of nodes is given by $N=k^{d_{I}}$ with $k$ the number of evaluation points per dimension and $d_{I}$ the integral dimensionality. Due to this exponential scaling, dimensionality reduction strategies are necessary. Note that while the second transformation eliminated the non-polynomial Copula density $c(\bm{u})$, improving numerical accuracy, there is still no exactness guarantee. Gauss-Hermite quadrature is exact for polynomials up to degree 2k12k-1 weighted by a normal PDF, but $g(\bm{F}^{-1}(\bm{\Phi}(\bm{Lz})))$ does not necessarily behave like a polynomial of degree $\leq 2k-1$ in Gaussian latent space.
@@ II-C1 Numerical Improvements
The dimensionality-reduction strategy used to make quadrature integration tractable stems from two key observations: First, as noted in [10], inner products between polynomials can be expanded to a sum of monomial expectation values. For instance, let $P_{m}(\bm{\xi}),P_{n}(\bm{\xi})$ be two arbitrary polynomials constructed from the set of multivariate monomials of degree $\leq\nu$ $\{1,\xi_{1},\ldots,\xi_{d}^{\nu}\}$. Then the inner product
$\left\langle P_{m},P_{n}\right\rangle$ $=d_{0}\!\left\langle 1,1\right\rangle+d_{1}\!\left\langle 1,\xi_{1}\right\rangle+\ldots+d_{D}\!\left\langle\xi_{d}^{\nu},\xi_{d}^{\nu}\right\rangle$
@@
where $d_{i},e_{i}$ are the appropriate coefficients. Second, the expectation values of monomials only depend on dimensions appearing in the monomial, for instance
@@
where $p_{1,2}(\xi_{1},\xi_{2})$ is a reduced joint PDF. For elliptic Copulas they have identical structure to the full joint PDF defined in (7), just with less input variables. For other Copulas, the reduced joint PDFS might need to be retrieved separately.
Now, instead of directly applying an orthonormalization scheme, the monomial expectation values of all pairs of monomials in the chosen set can be evaluated beforehand. When applying the orthonormalization scheme, the inner products can be expanded as in (18) and the expectation values substituted.
Crucially, precomputing expectation values can be done in a parallel fashion and naturally limits the maximum integral dimension to be evaluated to $2\nu$. This because the highest dimensional expectation values to be evaluated are of the form $\mathbb{E}[\xi_{1}\xi_{2}\ldots\xi_{2\nu}]$. This approach decouples the integral dimension $d_{I}$ from the random variable dimension $d$.
In the case where multiple random functions $q_{i}(\bm{\xi_{i}})$ only depending on subsets $\bm{\xi_{i}}$ of the random vector are expanded to the same polynomial basis, a similar approach can be used for the expansion coefficients. Here the expectation values $\mathbb{E}[q_{i}(\bm{\xi_{i}})],\ldots,\mathbb{E}[q_{i}(\bm{\xi_{i}})\xi_{d}^{\nu}]$ can be computed before with reduced dimension. Note, that the reduction in this case is dependent on the dimensionality of the random functions $q_{i}(\bm{\xi_{i}})$.
Finally, the set of monomials from which the polynomial basis is constructed can be tuned. For instance, assume the random functions $q_{i}(\bm{\xi_{i}})$ are models of PV power output at locations $i$ with $\bm{\xi_{i}}$ the local weather conditions like irradiation and temperature. Physically meaningless monomials such as the product of squares of the irradiations at two different locations can be dropped from the set without significantly affecting the expansion accuracy.